\begin{thebibliography}{99}


\bibitem{toptev}  F.~Abe {\it et al.}  [CDF Collaboration],
  %``Observation of top quark production in $\bar{p}p$ collisions,''
  Phys.\ Rev.\ Lett.\  {\bf 74} (1995) 2626;
  %%CITATION = PRLTA,74,2626;%%
 S.~Abachi {\it et al.}  [D0 Collaboration],
  %``Observation of the top quark,''
  Phys.\ Rev.\ Lett.\  {\bf 74},  (1995) 263.
  %%CITATION = PRLTA,74,2632;%%

%\cite{Nason:1987xz}
\bibitem{Nason:1987xz}
  P.~Nason, S.~Dawson and R.K.~Ellis,
  %``The Total Cross-Section for the Production of Heavy Quarks in Hadronic
  %Collisions,''
  Nucl.\ Phys.\  B {\bf 303} (1988)  607;
  %%CITATION = NUPHA,B303,607;%%
%
%\cite{Nason:1989zy}
%\bibitem{Nason:1989zy}
%  P.~Nason, S.~Dawson and R.K.~Ellis,
  %``The One Particle Inclusive Differential Cross-Section for Heavy Quark
  %Production in Hadronic Collisions,''
  Nucl.\ Phys.\  B {\bf 327} (1989)  49
  [Erratum-ibid.\  B {\bf 335} (1990) 260].
  %%CITATION = NUPHA,B327,49;%%
%
%\cite{Beenakker:1988bq}
%\bibitem{Beenakker:1988bq}
  W.~Beenakker {\it et al.},
  %``QCD Corrections to Heavy Quark Production in p anti-p Collisions,''
  Phys.\ Rev.\  D {\bf 40} (1989)  54.
  %%CITATION = PHRVA,D40,54;%%
%
%\cite{Beenakker:1990maa}
%\bibitem{Beenakker:1990maa}
  W.~Beenakker {\it et al.},
  %``QCD corrections to heavy quark production in hadron hadron collisions,''
  Nucl.\ Phys.\  B {\bf 351} (1991)  507.
  %%CITATION = NUPHA,B351,507;%%
%
%\cite{Mangano:1991jk}
%\bibitem{Mangano:1991jk}
  M.L.~Mangano, P.~Nason and G.~Ridolfi,
  %``Heavy quark correlations in hadron collisions at next-to-leading order,''
  Nucl.\ Phys.\  B {\bf 373} (1992)  295.
  %%CITATION = NUPHA,B373,295;%%
%\cite{Korner:2002hy}
%
%\bibitem{Korner:2002hy}
  J.G.~K\"orner and Z.~Merebashvili,
  %``One-loop corrections to four-point functions with two external massive
  %fermions and two external massless partons,''
  Phys.\ Rev.\  D {\bf 66} (2002)  054023.
  %%CITATION = PHRVA,D66,054023;%%
%
%\cite{Bernreuther:2004jv}
%\bibitem{Bernreuther:2004jv}
  W.~Bernreuther {\it et al.},
  %``Top quark pair production and decay at hadron colliders,''
  Nucl.\ Phys.\  B {\bf 690} (2004) 81.
  %%CITATION = NUPHA,B690,81;%%


\bibitem{Kidonakis:1997gm}
  N.~Kidonakis and G.~Sterman,
  %``Resummation for QCD hard scattering,''
  Nucl.\ Phys.\  B {\bf 505} (1997) 321.
  %%CITATION = NUPHA,B505,321;%%
%
%\bibitem{Bonciani:1998vc}
  R.~Bonciani {\it et al.},
  %``NLL resummation of the heavy-quark hadroproduction cross-section,''
  Nucl.\ Phys.\  B {\bf 529} (1998) 424.
  %%CITATION = NUPHA,B529,424;%%
%
%\bibitem{Cacciari:2003fi}
  M.~Cacciari {\it et al.},
  %``The t anti-t cross-section at 1.8-TeV and 1.96-TeV: A study of the
  %systematics due to parton densities and scale dependence,''
  JHEP {\bf 0404} (2004) 068.
  %%CITATION = JHEPA,0404,068;%%

%%%%%% NLO corrections to top quark pair production

  %\cite{Kuhn:2005it}
\bibitem{Kuhn:2005it}
  J.H.~K\"uhn, A.~Scharf and P.~Uwer,
  %``Electroweak corrections to top-quark pair production in quark-antiquark
  %annihilation,''
  Eur.\ Phys.\ J.\  C {\bf 45} (2006) 139;
  Eur.\ Phys.\ J.\  C {\bf 51} (2007) 37.
  %%CITATION = EPHJA,C45,139;%%
  %%CITATION = EPHJA,C51,37;%%
%
  %\cite{Bernreuther:2006vg}
%\bibitem{Bernreuther:2006vg}
  W.~Bernreuther, M.~F\"ucker and Z.G.~Si,
  %``Weak interaction corrections to hadronic top quark pair production,''
  Phys.\ Rev.\  D {\bf 74} (2006) 113005;  arXiv:0804.1237.
  %%CITATION = ARXIV:0804.1237;%%
  %%CITATION = PHRVA,D74,113005;%%
%
%\bibitem{Beenakker:1993yr}
  W.~Beenakker {\it et al.},
  %``Electroweak one loop contributions to top pair production in hadron
  %colliders,''
  Nucl.\ Phys.\  B {\bf 411} (1994) 343.
  %%CITATION = NUPHA,B411,343;%%


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%



\bibitem{Moch:2008qy}
  S.~Moch and P.~Uwer,
  %``Theoretical status and prospects for top-quark pair production at hadron
  %colliders,''
  Phys.\ Rev.\  D {\bf 78} (2008) 034003.
  %%CITATION = PHRVA,D78,034003;%%
%
%\bibitem{Cacciari:2008zb}
  M.~Cacciari {\it et al.},
  %``Updated predictions for the total production cross sections of top and of
  %heavier quark pairs at the Tevatron and at the LHC,''
  JHEP {\bf 0809} (2008) 127.
  %%CITATION = ARXIV:0804.2800;%%
%
%\bibitem{Kidonakis:2008mu}
  N.~Kidonakis and R.~Vogt,
  %``The theoretical top quark cross section at the Tevatron and the LHC,''
  arXiv:0805.3844.
  %%CITATION = ARXIV:0805.3844;%%

%\cite{Dittmaier:2007wz}
\bibitem{Dittmaier:2007wz}
  S.~Dittmaier, P.~Uwer and S.~Weinzierl,
  %``NLO QCD corrections to t anti-t + jet production at hadron colliders,''
  Phys.\ Rev.\ Lett.\  {\bf 98} (2007) 262002;
  %%CITATION = PRLTA,98,262002;%%
%\cite{Dittmaier:2008uj}
%\bibitem{Dittmaier:2008uj}
%  S.~Dittmaier, P.~Uwer and S.~Weinzierl,
  %``Hadronic top-quark pair production in association with a hard jet at
  %next-to-leading order QCD: Phenomenological studies for the Tevatron and the
  %LHC,''
  arXiv:0810.0452.
  %%CITATION = ARXIV:0810.0452;%%



\bibitem{Czakon:2007ej}
  M.~Czakon, A.~Mitov and S.~Moch,
  %``Heavy-quark production in massless quark scattering at two loops in QCD,''
  Phys.\ Lett.\  B {\bf 651} (2007) 147;
  %%CITATION = PHLTA,B651,147;%%
%
%\cite{Czakon:2007wk}
%\bibitem{Czakon:2007wk}
%  M.~Czakon, A.~Mitov and S.~Moch,
  %``Heavy-quark production in gluon fusion at two loops in QCD,''
  Nucl.\ Phys.\  B {\bf 798} (2008) 210.
  %%CITATION = NUPHA,B798,210;%%


\bibitem{Anastasiou:2000kg}
  Z.~Bern, L.J.~Dixon and A.~Ghinculov,
  %``Two-loop correction to Bhabha scattering,''
  Phys.\ Rev.\  D {\bf 63} (2001) 053007;
  %%CITATION = PHRVA,D63,053007;%%
 C.~Anastasiou {\it et al.},
  %``Two-loop QCD corrections to q anti-q --> q' anti-q',''
  Nucl.\ Phys.\  B {\bf 601} (2001) 318;
  %%CITATION = NUPHA,B601,318;%%
 E.W.N.~Glover,
  %``Two-loop QCD helicity amplitudes for massless quark quark scattering,''
  JHEP {\bf 0404} (2004) 021.
  %%CITATION = JHEPA,0404,021;%%

\bibitem{Anastasiou:2001sv}
  %  C.~Anastasiou {},
  C.~Anastasiou {\it et al.}
  %``Two-loop QCD corrections to massless quark gluon scattering,''
  Nucl.\ Phys.\  B {\bf 605} (2001) 486;
  %%CITATION = NUPHA,B605,486;%%
  E.W.N.~Glover and M.E.~Tejeda-Yeomans,
  %``Two-loop QCD helicity amplitudes for massless quark massless gauge  boson
  %scattering,''
  JHEP {\bf 0306} (2003) 033;
  %%CITATION = JHEPA,0306,033;%%
  Z.~Bern, A.~De Freitas and L.J.~Dixon,
  %``Two-loop helicity amplitudes for quark gluon scattering in QCD and  gluino
  %gluon scattering in supersymmetric Yang-Mills theory,''
  JHEP {\bf 0306} (2003) 028.
  %%CITATION = JHEPA,0306,028;%%


  %\cite{Czakon:2008zk}
\bibitem{Czakon:2008zk}
  M.~Czakon,
  %``Tops from Light Quarks: Full Mass Dependence at Two-Loops in QCD,''
  Phys.\ Lett.\  B {\bf 664} (2008) 307.
  %%CITATION = ARXIV:0803.1400;%%

\bibitem{US}
  R.~Bonciani {\it et al.},
  %``Two-Loop Fermionic Corrections to Heavy-Quark Pair Production: the
  %Quark-Antiquark Channel,''
  JHEP {\bf 0807} (2008) 129.
  %%CITATION = JHEPA,0807,129;%%

  %\cite{Korner:2005rg}
\bibitem{Korner:2005rg}
  J.G.~K\"orner, Z.~Merebashvili and M.~Rogal,
  %``One-loop amplitudes for four-point functions with two external massive
  %quarks and two external massless partons up to O(epsilon**2),''
  Phys.\ Rev.\  D {\bf 73}  (2006)  034030;
  %J.~G.~Korner, Z.~Merebashvili and M.~Rogal,
  %``NNLO ${\cal O}(\alpha_s^4)$ results for heavy quark pair production in
  %quark--antiquark collisions: The one-loop squared contributions,''
  Phys.\ Rev.\  D {\bf 77} (2008) 094011.
  %%CITATION = PHRVA,D77,094011;%%
   C.~Anastasiou and S.~M.~Aybat,
 %``The one-loop gluon amplitude for heavy-quark production at NNLO,''
 arXiv:0809.1355;
  B.~Kniehl {\it et al.},
 %``Heavy quark pair production in gluon fusion at next-to-next-to-leading
 %${\cal O}(\alpha_s^4)$ order: One-loop squared contributions,''
 arXiv:0809.3980.
 %%CITATION = ARXIV:0809.3980;%%
 %%CITATION = ARXIV:0809.1355;%%
  %%CITATION = PHRVA,D73,034030;%%

 \bibitem{qgraf}
 P.~Nogueira,
  %``Automatic Feynman graph generation,''
  J.\ Comput.\ Phys.\  {\bf 105} (1993) 279.
  %%CITATION = JCTPA,105,279;%%

\bibitem{FORM} J.A.M.\ Vermaseren, Symbolic Manipulation with
{\tt FORM}, Version 2, CAN, Amsterdam, 1991; math-th/0010025.
%%CITATION = MATH-TH 0010025;%%

\bibitem{electronloop}
  R.~Bonciani {\it et al.},
  %``Planar box diagram for the (N(F) = 1) 2-loop QED virtual corrections to
  %Bhabha scattering,''
  Nucl.\ Phys.\  B {\bf 681} (2004) 261
  [Erratum-ibid.\  B {\bf 702} (2004) 364].
  %%CITATION = NUPHA,B681,261;%%

\bibitem{hfbha}
  R.~Bonciani, A.~Ferroglia and A.A.~Penin,
  %``Two-Loop Heavy-Flavor Contribution to Bhabha Scattering,''
  Phys.\ Rev.\ Lett.\ {\bf 100} (2008) 131601;
  %%CITATION = PRLTA,100,131601;%%
  %\bibitem{hfbhaJHEP}
  %  R.~Bonciani, A.~Ferroglia and A.~A.~Penin,
  %``Calculation of the Two-Loop Heavy-Flavor Contribution to Bhabha
  %Scattering,''
  JHEP {\bf 0802} (2008) 080;
  %%CITATION = JHEPA,0802,080;%%
 S.~Actis {\it et al.},
  %``Two-Loop Fermionic Corrections to Massive Bhabha Scattering,''
  Nucl.\ Phys.\  B {\bf 786} (2007) 26;
  %%CITATION = NUPHA,B786,26;%%
  %``Virtual Hadronic and Leptonic Contributions to Bhabha Scattering,''
  Phys.\ Rev.\ Lett.\  {\bf 100} (2008) 131602.
  %%CITATION = PRLTA,100,131602;%%


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% LAPORTA ALGORITHM

\bibitem{Laportaalgorithm}
  S.~Laporta and E.~Remiddi,
  %``The analytical value of the electron (g-2) at order alpha~3 in QED,''
  Phys.\ Lett.\ B {\bf 379} (1996) 283;
  %%CITATION = HEP-PH 9602417;%%
  S.~Laporta,
  %``High-precision calculation of multi-loop Feynman integrals by  difference
  %equations,''
  Int.\ J.\ Mod.\ Phys.\ A {\bf 15} (2000) 5087.
  %%CITATION = HEP-PH 0102033;%%
  %
  %\bibitem{Tka}
  F.V.~Tkachov,
  %``A Theorem On Analytical Calculability Of Four Loop Renormalization Group
  %Functions,''
  Phys.\ Lett.\ B {\bf 100} (1981) 65;
  %%CITATION = PHLTA,B100,65;%%
  K.G.~Chetyrkin and F.V.~Tkachov,
  %``Integration By Parts: The Algorithm To Calculate Beta Functions In 4 Loops,''
  Nucl.\ Phys.\ B {\bf 192} (1981) 159.
  %%CITATION = NUPHA,B192,159;%%
%\bibitem{Anastasiou:2004vj}
  C.~Anastasiou and A.~Lazopoulos,
  %``Automatic integral reduction for higher order perturbative  calculations,''
  JHEP {\bf 0407} (2004)  046.
  %%CITATION = JHEPA,0407,046;%%


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% WHERE TO FIND THE MIs

%\cite{Fleischer:1999hp}
\bibitem{Fleischer:1999hp}
  J.~Fleischer, M.~Y.~Kalmykov and A.~V.~Kotikov,
  %``Two-loop self-energy master integrals on shell,''
  Phys.\ Lett.\  B {\bf 462} (1999) 169.
  %%CITATION = PHLTA,B462,169;%%
%
%\bibitem{Argeri:2002wz}
  M.~Argeri, P.~Mastrolia and E.~Remiddi,
  %``The analytic value of the sunrise self-mass with two equal masses and  the
  %external invariant equal to the third squared mass,''
  Nucl.\ Phys.\  B {\bf 631} (2002)  388.
  %%CITATION = NUPHA,B631,388;%%
%
%\bibitem{Bonciani:2003te}
  R.~Bonciani, P.~Mastrolia and E.~Remiddi,
  %``Vertex diagrams for the QED form factors at the 2-loop level,''
  Nucl.\ Phys.\  B {\bf 661} (2003)  289
  [Erratum-ibid.\  B {\bf 702} (2004)  359].
  %%CITATION = NUPHA,B661,289;%%
%
%\bibitem{Bonciani:2003hc}
%  R.~Bonciani, P.~Mastrolia and E.~Remiddi,
  %``Master integrals for the 2-loop QCD virtual corrections to the
  %forward-backward asymmetry,''
  Nucl.\ Phys.\  B {\bf 690} (2004)  138.
  %%CITATION = NUPHA,B690,138;%%
%
%\bibitem{Aglietti:2003yc}
  J.~Fleischer, A.V.~Kotikov and O.L.~Veretin,
  %``Analytic two-loop results for selfenergy- and vertex-type diagrams  with
  %one non-zero mass,''
  Nucl.\ Phys.\  B {\bf 547} (1999) 343;
  %%CITATION = NUPHA,B547,343;%%
  U.~Aglietti and R.~Bonciani,
  %``Master integrals with one massive propagator for the two-loop  electroweak
  %form factor,''
  Nucl.\ Phys.\  B {\bf 668} (2003)  3;
  %%CITATION = NUPHA,B668,3;%%
%  U.~Aglietti and R.~Bonciani,
  %``Master integrals with 2 and 3 massive propagators for the 2-loop
  %electroweak form factor: Planar case,''
  Nucl.\ Phys.\  B {\bf 698} (2004) 277.
  %%CITATION = NUPHA,B698,277;%%
%
%\bibitem{DK}
  A.I.~Davydychev and M.Y.~Kalmykov,
  %``Massive Feynman diagrams and inverse binomial sums,''
  Nucl.\ Phys.\  B {\bf 699} (2004) 3.
  %%CITATION = NUPHA,B699,3;%%
%
%\bibitem{CGR}
  M.~Czakon, J.~Gluza and T.~Riemann,
  %``Master integrals for massive two-loop Bhabha scattering in QED,''
  Phys.\ Rev.\  D {\bf 71} (2005) 073009.
  %%CITATION = PHRVA,D71,073009;%%
%
%\bibitem{Bell:2006tz}
  G.~Bell,
  %``Higher order QCD corrections in exclusive charmless B decays,''
  arXiv:0705.3133.
  %%CITATION = ARXIV:0705.3133;%%





\bibitem{Bernreuther:2004ih}
  W.~Bernreuther, {\it et al.},
  %``Two-loop QCD corrections to the heavy quark form factors: The vector
  %contributions,''
  Nucl.\ Phys.\  B {\bf 706} (2005)  245;
  %%CITATION = NUPHA,B706,245;%%
  %``Two-loop QCD corrections to the heavy quark form factors: Axial vector
  %contributions,''
  Nucl.\ Phys.\  B {\bf 712} (2005) 229;
  %%CITATION = NUPHA,B712,229;%%
  %``Two-loop QCD corrections to the heavy quark form factors: Anomaly
  %contributions,''
  Nucl.\ Phys.\  B {\bf 723} (2005) 91;
  %%CITATION = NUPHA,B723,91;%%
  %``Decays of scalar and pseudoscalar Higgs bosons into fermions: Two-loop  QCD
  %corrections to the Higgs-quark-antiquark amplitude,''
  Phys.\ Rev.\  D {\bf 72} (2005) 096002;
  %%CITATION = PHRVA,D72,096002;%%
 %``QCD corrections to static heavy quark form factors,''
  Phys.\ Rev.\ Lett.\  {\bf 95} (2005) 261802.
  %%CITATION = PRLTA,95,261802;%%


\bibitem{bhabha}
 R.~Bonciani and A.~Ferroglia,
  %``Two-loop Bhabha scattering in QED,''
  Phys.\ Rev.\  D {\bf 72} (2005) 056004.
  %%CITATION = PHRVA,D72,056004;%%


%%%%%%%%%%%%%%%% DIFFERENTIAL EQUATIONS

\bibitem{DiffEq}
  A.V.~Kotikov,
  %``Differential Equations Method: New Technique For Massive Feynman Diagrams
  %Calculation,''
  Phys.\ Lett.\ B {\bf 254} (1991) 158;
  %%CITATION = PHLTA,B254,158;%%
  %A.~V.~Kotikov,
  %``Differential equations method: The Calculation of vertex type Feynman
  %diagrams,''
  Phys.\ Lett.\ B {\bf 259} (1991) 314;
  %%CITATION = PHLTA,B259,314;%%
  %A.~V.~Kotikov,
  %``Differential equation method: The Calculation of N point Feynman diagrams,''
  Phys.\ Lett.\ B {\bf 267} (1991) 123;
  %%CITATION = PHLTA,B267,123;%%
  E.~Remiddi,
  %``Differential equations for Feynman graph amplitudes,''
  Nuovo Cim.\ A {\bf 110} (1997) 1435;
  %%CITATION = HEP-TH 9711188;%%
  T.~Gehrmann and E.~Remiddi,
  %``Differential equations for two-loop four-point functions,''
  Nucl.\ Phys.\  B {\bf 580} (2000) 485;
  %%CITATION = NUPHA,B580,485;%%
  M.~Argeri and P.~Mastrolia,
  %``Feynman Diagrams and Differential Equations,''
  Int.\ J.\ Mod.\ Phys.\  A {\bf 22} (2007) 4375.
  %%CITATION = IMPAE,A22,4375;%%


%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% HPLs


\bibitem{HPLs}
  E.~Remiddi and J.A.M.~Vermaseren,
  %{\it ``Harmonic polylogarithms,''}
  Int.\ J.\ Mod.\ Phys.\ A {\bf 15} (2000) 725;
  %%CITATION = HEP-PH 9905237;%%
  %\bibitem{NUMHPL}
  T.~Gehrmann and E.~Remiddi,
  %{\it ``Numerical evaluation of harmonic polylogarithms,'' }
  Comput.\ Phys.\ Commun.\  {\bf 141} (2001) 296;
  %%CITATION = CPHCB,141,296;%%
  %T.~Gehrmann and E.~Remiddi,
  %{\it ``Numerical evaluation of two-dimensional harmonic polylogarithms,'' }
  Comput.\ Phys.\ Commun.\  {\bf 144} (2002) 200.
  %%CITATION = CPHCB,144,200;%%
  J.~Vollinga and S.~Weinzierl,
  %``Numerical evaluation of multiple polylogarithms,''
  Comput.\ Phys.\ Commun.\  {\bf 167} (2005) 177;
  %%CITATION = CPHCB,167,177;%%
  %\bibitem{Daniel1}
  D.~Ma\^{i}tre,
  %{\it ``HPL, a mathematica implementation of the harmonic polylogarithms'',}
  Comput.\ Phys.\ Commun.\ {\bf 174} (2006) 222;
  %%CITATION = CPHCB,174,222;%%
  %``Extension of HPL to complex arguments,''
  hep-ph/0703052.
  %%CITATION = HEP-PH/0703052;%%


\end{thebibliography}

\documentclass[slac_one]{revtex4}
\usepackage{graphicx}
\usepackage{fancyhdr}
\begin{document}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bfm}[1]{\mbox{\boldmath$#1$}}
\newcommand{\bff}[1]{\mbox{\scriptsize\boldmath${#1}$}}
\newcommand{\al}{\alpha}
\newcommand{\bt}{\beta}
\newcommand{\lm}{\lambda}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\gm}{\gamma}
\newcommand{\Gm}{\Gamma}
\newcommand{\dl}{\delta}
\newcommand{\Dl}{\Delta}
\newcommand{\ep}{\varepsilon}
\newcommand{\vep}{\varepsilon}
\newcommand{\kp}{\kappa}
\newcommand{\Lm}{\Lambda}
\newcommand{\om}{\omega}
\newcommand{\pa}{\partial}
\newcommand{\nn}{\nonumber}
\newcommand{\dd}{\mbox{d}}
\newcommand{\grtsim}{\mbox{\raisebox{-3pt}{$\stackrel{>}{\sim}$}}}
\newcommand{\lessim}{\mbox{\raisebox{-3pt}{$\stackrel{<}{\sim}$}}}
\newcommand{\uk}{\underline{k}}
\newcommand{\gsim}{\;\rlap{\lower 3.5 pt \hbox{$\mathchar \sim$}} \raise 1pt \hbox {$>$}\;}
\newcommand{\lsim}{\;\rlap{\lower 3.5 pt \hbox{$\mathchar \sim$}} \raise 1pt \hbox {$<$}\;}
\newcommand{\Li}{\mbox{Li}}
\newcommand{\bc}{\begin{center}}
\newcommand{\ec}{\end{center}}

\def\lapprox{\lower .7ex\hbox{$\;\stackrel{\textstyle <}{\sim}\;$}}
\def\gapprox{\lower .7ex\hbox{$\;\stackrel{\textstyle >}{\sim}\;$}}

\title{Fermionic Corrections to the Heavy-Quark Pair Production in the Quark-Antiquark Channel}

\author{R.~Bonciani, A.~Ferroglia, T.~Gehrmann, C.~Studerus}
\affiliation{University of Z\"urich, Z\"urich, CH 8057, Switzerland }
\author{D.~Ma\^itre}
\affiliation{SLAC, Stanford, CA 94025, USA}

\begin{abstract}
We describe the analytic calculation of the fermionic two-loop QCD
corrections to the heavy-quark pair production process in the
quark-antiquark channel.
\end{abstract}

\maketitle

\thispagestyle{fancy}

\section{INTRODUCTION}

The top quark is the heaviest fermion of the Standard Model. Since its discovery at
the Fermilab Tevatron~\cite{toptev}, its mass has  been measured to within a few
percent, while its production cross  section and couplings are currently known with
larger uncertainty.  With the large number of top quarks expected to be produced at
the  LHC, the study of its properties will become precision physics. To  interpret
these upcoming precision data, equally precise theoretical  predictions are
mandatory. These demand foremost the calculation  of higher order corrections in
perturbative QCD.

At present, the top quark pair production cross section is known to next-to-leading
order (NLO) in the QCD coupling constant~\cite{Nason:1987xz}. For this process, the
resummation of next-to-leading logarithmically enhanced  corrections (NLL) improves
upon the fixed-order NLO prediction~\cite{Kidonakis:1997gm}. Electroweak one-loop
corrections to $t\bar t$ production are equally available \cite{Kuhn:2005it}.
For the top quark pair production cross section, which is expected to be measured to
within a few percent accuracy, the currently available theoretical prediction is not sufficiently
precise. Recent studies~\cite{Moch:2008qy} indicate a scale uncertainty on these
predictions of 7\%, and a parton distribution uncertainty of 6\%. While the latter
may be improved upon by more precise determinations of the parton distribution
functions at HERA and LHC, the former requires the calculation of perturbative
corrections at next-to-next-to-leading order (NNLO) in QCD.

The calculation of the full NNLO corrections to the top quark pair production cross
section requires three types of ingredients: two-loop matrix elements for $q\bar q
\to t\bar t$ and $gg\to t\bar t$, one-loop matrix elements for hadronic production
of $t\bar t+$(1 parton)  and tree-level matrix elements for hadronic production of
$t\bar t+$(2 partons). The latter two ingredients were computed previously in the
context of the NLO corrections to $t\bar t$+jet production~\cite{Dittmaier:2007wz}.
They contribute to the  $t\bar t$ production cross section through configurations
where up to two final state partons can be unresolved (collinear or soft), and their
implementation thus may require further developments of subtraction techniques at
NNLO.

Both two-loop matrix elements were computed analytically in the small-mass expansion
limit $s, |t|, |u| \gg m^2$ in~\cite{Czakon:2007ej}, starting from the previously
known massless two-loop matrix elements for $q\bar q\to q'\bar
q'$~\cite{Anastasiou:2000kg} and $gg\to q\bar q$~\cite{Anastasiou:2001sv}. An exact
numerical representation of the two-loop matrix element  $q\bar q \to t\bar t$ has
been obtained very recently~\cite{Czakon:2008zk}. In \cite{US} we computed all
two-loop contributions to $q\bar q \to t\bar t$ arising from closed fermion loops in
a compact analytic form, providing also a first independent validation of the
results of~\cite{Czakon:2007ej,Czakon:2008zk}. Our results allow for a fast
numerical evaluation and permit the analytical study of the cross section near
threshold. In the rest of this proceeding we
briefly discuss the structure of the two-loop fermionic corrections and
the calculational techniques employed to evaluate them.
%in \cite{US}.

\section{STRUCTURE}

The scattering process we consider is $q(p_1) + \overline{q}(p_2)
\to  t(p_3) + \overline{t}(p_4)$
in Euclidean kinematics, where $p_i^2 = 0$ for $i=1,2$  and $p_j^2
= -m^2$ for  $i=3,4$. The Mandelstam variables are defined as
follows:
$s = -\left(p_1 + p_2 \right)^2$, $t = -\left(p_1 - p_3
\right)^2$, $u = -\left(p_1 - p_4 \right)^2$.  Conservation of
momentum implies that $s +t +u = 2 m^2$.
The squared matrix element (averaged over the spin and color of
the incoming quarks and summed over the spin of the outgoing
ones), calculated in $d = 4 -2 \varepsilon$ dimensions, can be
expanded in powers of the strong coupling constant $\alpha_S$ as
follows:
\be \label{M2} |\mathcal{M}|^2(s,t,m,\varepsilon) = \frac{4 \pi^2
\alpha_S^2}{N_c^2} \left[{\mathcal A}_0 + \left(\frac{\alpha_s}{
\pi} \right) {\mathcal A}_1 + \left(\frac{\alpha_s}{ \pi}
\right)^2 {\mathcal A}_2 + {\mathcal O}\left(
\alpha_s^3\right)\right] \, . \ee
The tree-level amplitude involves a single diagram
 and its contribution to Eq.~(\ref{M2}) is
given by
\be
{\mathcal A}_0=  4  N_c \, C_F \left[
\frac{(t-m^2)^2+(u-m^2)^2}{s^2} + \frac{2 m^2}{s} -
\varepsilon \right] \, ,
\ee
where  $N_c$ is the number of colors and  $C_F = (N_c^2-1)/2N_c$.

The NLO term ${\mathcal A}_1$  in Eq.~(\ref{M2}) arises from the interference of
one-loop diagrams with the tree-level amplitude \cite{Nason:1987xz}.
The NNLO term ${\mathcal A}_2$ consists of two parts, the interference of two-loop
diagrams with the Born amplitude and the interference of one-loop diagrams among
themselves: $ {\mathcal A}_2 = {\mathcal A}_2^{(2\times 0)} + {\mathcal
A}_2^{(1\times 1)}$. The latter term ${\mathcal A}_2^{(1\times 1)}$ was studied
extensively  in \cite{Korner:2005rg}. ${\mathcal A}_2^{(2\times 0)}$
%originating
%from the two-loop  diagrams,
can be decomposed according to color and flavor
structures as follows:
\bea
{\mathcal A}_2^{(2\times 0)}  &=&  N_c C_F \left[ N_c^2 A
+B +\frac{C}{N_c^2}  + N_l \left( N_c D_l + \frac{E_l}{N_c}
\right)
+ N_h \left( N_c D_l + \frac{E_l}{N_c} \right)
+ N_l^2 F_l + N_l N_h F_{lh} + N_h^2 F_h
\right] \, ,
\label{colstruc}
\eea
where $N_l$ and $N_h$ are the number of light- and heavy-quark flavors,
respectively. The coefficients $A,B,\ldots,F_h$ in Eq.~(\ref{colstruc}) are
functions of $s$, $t$, $m$, and
$\varepsilon$. These quantities were calculated in  \cite{Czakon:2007ej} in the
approximation $s,|t|,|u| \gg m^2$. For a fully differential description of top quark
pair production at NNLO, the complete mass dependence of  ${\mathcal A}_2^{(2\times
0)}$ is required. An exact numerical expression for it has been obtained in
\cite{Czakon:2008zk}. In \cite{US}, we derived exact analytic expressions for all
the terms in Eq.~(\ref{colstruc}) arising from two-loop diagrams involving at least
a fermion loop (i.e. the coefficients $D_i,E_i,F_j$ with $i=l,h$ and $j=l,h,lh$),
providing also an independent confirmation of the  results of
\cite{Czakon:2007ej,Czakon:2008zk}.

\section{CALCULATION\label{sec:calc}}

The two-loop Feynman diagrams for $q\bar q\to t\bar t$ were generated with
QGRAF~\cite{qgraf}. The interference with the tree-level amplitude, as well as the
color and Dirac algebra, were simplified by using a FORM~\cite{FORM} code. Out of
the $\sim$ 200 two-loop diagrams contributing to the amplitude, about 60 are
proportional to $N_l$ and/or $N_h$. There is only one two-loop box topology
contributing to the $N_l$ part of the squared amplitude, and a single other two-loop
box topology proportional to $N_h$. These two box topologies are very similar to the
ones encountered in the evaluation of the two-loop QED corrections to Bhabha
scattering \cite{electronloop,hfbha}, and can be evaluated with the same techniques.

All two-loop integrals appearing in these amplitudes are reduced to a set of master
integrals (MIs) using two independent implementations of the Laporta algorithm
\cite{Laportaalgorithm}.
Only part of these MIs were available in the literature~\cite{Fleischer:1999hp}
from previous two-loop calculations of the heavy quark form factors
\cite{Bernreuther:2004ih} and amplitudes for Bhabha scattering
\cite{electronloop,hfbha,bhabha}. The remaining MIs
%, including four two-loop
%four legs integrals,
were evaluated in \cite{US} by employing the differential equation method
\cite{DiffEq}.

All the MIs were calculated in the non-physical region $s<0$.
The transcendental functions appearing in the MIs are one- and two-dimensional
harmonic polylogarithms (HPLs) \cite{HPLs} of maximum weight four and three,
respectively. Both sets of functions can be rewritten in  terms of conventional
Nielsen's polylogarithms.

Following the procedure outlined in the present section, it was possible to obtain
the expression of the bare squared matrix elements involving diagrams proportional
to $N_l$ and/or $N_h$. The UV divergencies were renormalized in a mixed scheme
described in detail in \cite{US}.
In order to cross check our analytical results, we expanded them in the $s,|t|, |u|
\gg m^2$ limit. The first term in the expansion agrees with the results published in
\cite{Czakon:2007ej}; the second order term agrees with the results found in the
{\tt Mathematica} files  included in the arXiv version of \cite{Czakon:2008zk}. We
also find complete agreement with the numerical result of Table~3 in
\cite{Czakon:2008zk}, corresponding to a phase  space point in which the $s,|t|, |u|
\gg m^2$ approximation cannot be applied.

\begin{acknowledgments}

Work supported by the Swiss National Science Foundation (SNF)
under contract 200020-117602. 
\end{acknowledgments}

\end{document}